# Code Difference Guided Fuzzing for FPGA Logic Synthesis Compilers via Bayesian Optimization

### v1

@@ IV-F Answer to RQ4
Approach. To evaluate the impact of the Bayesian selector on HDL code generation, we conducted an ablation study comparing the complexity and generation time of HDL codes produced with and without the Bayesian selector. Each scenario involved generating 100 HDL codes, each ranging from 700-1000 lines.
Results. As shown as Figure 10, the generation of HDL code without the Bayesian selector proved to be 7% faster on average compared to scenarios where the selector was used. However, the HDL codes generated without the Bayesian selector contained approximately 18% fewer blocks and connections, indicating reduced complexity. Furthermore, these codes exhibited 10% fewer reference relationships, suggesting that even though the Bayesian selector does not directly alter HDL code, its absence noticeably decreases the intricacy of the generated codes. Thus, incorporating the Bayesian selector enhances the complexity of HDL test cases without significantly impacting generation speed.
Conclusion. The Bayesian selector effectively increases the complexity of generated test cases without compromising the efficiency of the generation process.
@@ V Threats to Validity
Internal Threats. The primary threat to the internal validity of our method lies in the accuracy of LSC-Fuzz, as well as the reproduced models Verismith [7] and LegoHDL [8]. To mitigate errors in LSC-Fuzz, we meticulously double-checked our code. Additionally, we reproduced Verismith and LegoHDL by closely following the open-source code provided in their original publications, and we have thoroughly verified their accuracy. Furthermore, the model checking capabilities of Simulink HDL Coder can lead to increased processing times when generating large volumes of code. To address this, we optimized block placement to minimize model checking duration. Moving forward, we plan to intensify our testing of the FPGA logic Synthesis compiler by focusing on the generation of more complex test cases.
Figure 9: Blocks, connections and references generated by LSC-Fuzz, LegoHDL and Verismith
Figure 10: Blocks, connections and references comparison with Bayesian selector
External Threats. A significant threat to the external validity of LSC-Fuzz is the potential for triggering duplicate bugs during testing. To mitigate this, we have implemented strategies to reduce bug-triggering variants by excising blocks unrelated to the core issues, which enhances our ability to accurately understand and analyze bugs. Based on the developer’s verification, the bugs we reported are currently not duplicates, suggesting that our approach is effective in minimizing this threat.
@@ VI-A FPGA Logic Synthesis Testing
To overcome these limitations and enhance the thoroughness of FPGA logic synthesis compiler testing, Xu et al. introduced LegoHDL [8], which aims to ensure the accuracy and correctness of FPGA logic synthesis processes. LegoHDL employs HDLcoder to generate HDL code by constructing a CPS model. It extracts the AST from the HDL code and orchestrates the construction of the CPS model to facilitate interaction between CPS model construction and HDL code generation. Finally, LegoHDL utilizes Metamorphic testing to synthesize the HDL code using various synthesizers and verify the equivalence of the generated netlist files. However, LegoHDL still faces challenges with insufficient diversity, and its singular CPS construction approach does not adequately address deeper bugs.
@@ VI-B Metamorphic testing in Compiler testing
Metamorphic testing has also been applied in other fields. For example, SLEMI [25] proposed by Shaful et al. uses strategies such as deleting dead code to generate CPS variants to thoroughly test the Simulink compiler. Guo et al. proposed COMBAT [26], which uses the idea of live code mutation to generate CPS variants in order to more thoroughly test the Simulink compiler. In addition, the idea of Metamorphic testing has also been applied at the project level. Wu et al. proposed Mantra [27] to apply mutation testing to HDL and achieved good results.
@@ VII Conclusions and Future Work
In this paper, we proposed a novel method for FPGA logic synthesis compiler testing, named LSC-Fuzz. This method incorporates a Test-Program Generation component and a Bayesian selection component, enabling the generation of diverse HDL code for testing the FPGA logic synthesis compiler. Over a three-month period, our method successfully identified 16 bugs, 12 of which have been confirmed by developers. Going forward, we plan to explore how to integrate large models to guide the generation of even more diverse test cases, further enhancing our testing of FPGA logic synthesis compilers.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (No.62472062, No.62202079), the Dalian Excellent Young Project (2022RY35), the Fundamental Research Funds for the Central Universities (No.3132024257).

### v2

@@ IV-F Answer to RQ4
Figure 7: Blocks, Connections and Delay Comparison with Bayesian Selector and without Bayesian Selector
Approach. We conducted an ablation to quantify the effect of the Bayesian selector in SmootHDL. Two settings were evaluated: with selector and without selector. Each setting produced 100 HDL test cases within 700–1000 lines under the same tool chain and hardware configuration. We recorded generation time for each case. Structural metrics were extracted from the synthesized AST, which is block count and connection count. Timing complexity was measured as the critical-path delay reported by the synthesis timing analysis. All other parameters of the generator were fixed so that the selector was the only changing factor.
Results. The selector shifts the distributions of structure and timing upward with a modest increase in generation time. From the box-plots in Figure7, the median number of connections increases from about 175 without the selector to about 210 with the selector, an increase of roughly twenty percent; the inter quartile ranges show little overlap, indicating a consistent effect. The median number of blocks rises from about 31 to about 38, an increase of roughly twenty to twenty-five percent, again with clearly separated quartiles. The median critical-path delay increases from about 18 ns to about 20 ns, an increase of roughly ten to twelve percent; the distributions partly overlap, which suggests a moderate timing-complexity gain. Generation time shows a small upward shift, with the median moving from about 300 s to about 320–330 s, around seven to ten percent. Taken together, these observations indicate that the selector tends to guide the generator toward HDL designs with higher structural richness and somewhat deeper logic paths, while the additional runtime is limited under this experimental budget.
Conclusion. The Bayesian selector contributes to the effectiveness of SmootHDL by guiding the selection of variants with greater structural complexity and diversity, while maintaining comparable generation efficiency.
@@ V Threats to Validity
Internal Threats. The primary threat to the internal validity of our method lies in the accuracy of SmootHDL, as well as the reproduced models Verismith [9] and LegoHDL [7]. To mitigate errors in SmootHDL, we meticulously double-checked our code. Additionally, we reproduced Verismith and LegoHDL by closely following the open-source code provided in their original publications, and we have thoroughly verified their accuracy. Furthermore, the model checking capabilities of Simulink HDL Coder can lead to increased processing times when generating large volumes of code. To address this, we optimized block placement to minimize model checking duration. Moving forward, we plan to intensify our testing of the FPGA logic Synthesis compiler by focusing on the generation of more complex test cases.
External Threats. A significant threat to the external validity of SmootHDL is the potential for triggering duplicate bugs during testing. To mitigate this, we have implemented strategies to reduce bug-triggering variants by excising blocks unrelated to the core issues, which enhances our ability to accurately understand and analyze bugs. Based on the developers’s verification, the bugs we reported are currently not duplicates, suggesting that our approach is effective in minimizing this threat. Furthermore, the effectiveness of the Bayesian selector depends on the richness of the candidate pool and the granularity of the diversity features. When the available transformations yield candidates that are semantically similar at the FSM level (for example, simple pipelines or largely stateless datapaths), the posterior can concentrate early and the selector may provide limited benefit. In SmootHDL we mitigate this risk by enriching the pool through control-flow and timing oriented mutations, including state or path cloning, inhibited transitions, and scheduling variations. And we plan to study stopping rules that detect pool homogeneity and trigger automatic pool refresh, so that the selector remains effective when diversity declines.
@@ VI-A FPGA Logic Synthesis Testing
To overcome these limitations and enhance the thoroughness of FPGA logic synthesis compiler testing, Xu et al. introduced LegoHDL [7], which aims to ensure the accuracy and correctness of FPGA logic synthesis processes. LegoHDL employs HDL Coder to generate HDL code by constructing a Simulink Model. It extracts the AST from the HDL code and orchestrates the construction of the Simulink Model to facilitate interaction between Simulink Model construction and HDL code generation. Finally, LegoHDL utilizes Metamorphic testing to synthesize the HDL code using various synthesizers and verify the equivalence of the generated netlist files. However, LegoHDL still faces challenges with insufficient diversity, and its singular Simulink Model construction approach does not adequately address deeper bugs.
@@ VI-B Metamorphic testing in Compiler testing
Metamorphic testing has also been applied in other fields. For example, SLEMI [25] proposed by Shaful et al. uses strategies such as deleting dead code to generate Simulink Model variants to thoroughly test the Simulink compiler. Guo et al. proposed COMBAT [26], which uses the idea of live code mutation to generate Simulink Model variants in order to more thoroughly test the Simulink compiler. In addition, the idea of Metamorphic testing has also been applied at the project level. Wu et al. proposed Mantra [27] to apply mutation testing to HDL and achieved good results.
@@ VII Conclusions and Future Work
In this paper, we proposed a novel method for FPGA logic synthesis compiler testing, named SmootHDL. This method incorporates a Test-Program Generation component and a Bayesian selection component, enabling the generation of diverse HDL code for testing the FPGA logic synthesis compiler. Over a three-month period, our method successfully identified 16 bugs, 12 of which have been confirmed by developers. Going forward, we plan to explore how to integrate large models to guide the generation of even more diverse test cases, further enhancing our testing of FPGA logic synthesis compilers.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (No.62472062), the Dalian Excellent Young Project (2022RY35), the Fundamental Research Funds for the Central Universities (No.3132024257).